OpenStudy (anonymous):

If n-2 = 1/16, then n could be which of the following? A.-8 B.-4 C.-1/4 D.1/4 E.4

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

is that the only one

OpenStudy (anonymous):

I think I might go with B and E.

OpenStudy (anonymous):

um nope... n=33/16 doesn't it?

OpenStudy (anonymous):

Well, that's not one of your options, so I think we were assuming you meant n^-2.

OpenStudy (anonymous):

n^-2=1/16 n would be 4, E. Not B, -4..... -4^-2=-1/16

OpenStudy (anonymous):

1/n^2 = 1/16, n = plus or minus 4?

jhonyy9 (jhonyy9):

n-2=1/16 --- n=2+1/16 --- 16n=33 --- n=33/16